# OoDIS: Anomaly Instance Segmentation Benchmark

### v1

@@ 4 Evaluated Methods & Discussion of Results
[…]
Mask2Anomaly [32] is a model that uses auxilary data, but does not use an external model for instance segmentation. Common to other methods in the community [13, 35], the model uses auxiliary data from COCO [26] for guiding the anomaly scores that are grouped using connected components to form instance proposals. To reduce the number of false positives, Mask2Anomaly introduces a post-processing strategy. It computes the intersection with predicted in-distribution masks and uses class entropy to determine true instance proposals. The approach benefits from a powerful backbone and is effective in segmenting individual anomalous objects, however, it merges closely located anomalies (see Figure 3).
UGainS [29] is a method that uses both auxiliary data and an external generalist segmentation model, namely the segment anything model (SAM) [22]. The method uses the anomaly segmentation method RbA [28] based on Mask2Former [8], fine-tuned using data from COCO, to generate uncertainty regions. UGainS uses farthest point sampling to sample a number of points from these regions as prompts for SAM [22]. While the method produces accurate segmentation masks, it relies on two models to get predictions. A limited number of prompts leads to missed detections in smaller regions and increases the number of false positives in other areas. However, it demonstrates strong performance and produces well-separated instance masks.
(a) Ground Truth
(b) UGainS
(c) Mask2Anomaly
(d) U3HS
Figure 3: Qualitative comparison of the methods. The scene contains multiple grouped anomaly objects close to the camera and multiple smaller instances in the distance.
5 Conclusion
@@
Detecting and accurately segmenting anomaly instances on roads is a significant challenge, requiring an understanding of ’objectness’ without direct training on specific anomaly classes. In this work, we introduced a new benchmark for anomaly instance segmentation that integrates three popular anomaly datasets. The unified benchmark provides a diverse set of anomalies that vary in size, number of images, and annotation detail. We evaluate the performance of current methods for segmenting anomaly instances and provide intuition behind the results. Our results show that current techniques struggle particularly with distant and small objects, and with precise segmentation masks. The benchmark results suggest strong opportunities for advancement in the area. As autonomous vehicle technologies continue to evolve, driven by large amounts of data, it remains a challenge to capture all possible real-world situations. Our work addresses the need to evaluate instance segmentation as a step towards reliable autonomous driving.
@@ Acknowledgment
We acknowledge fruitful discussions with H. Blum, R. Chan, S. Gasperini and S. Rai, as well as a contribution of annotations for Fishyscapes by H. Blum, and help with the benchmark submission from S. Gasperini and S. Rai. M.R. an M.A. acknowledge support by the German Federal Ministry of Education and Research within the junior research group project “UnrEAL” (grant no. 01IS22069). Alexey Nekrasov’s research has been funded by BMBF project ”WestAI” (grant no. 01IS22094D). We thank M. Burdorf, G. Lydakis, C. Schmidt, and others in the lab for discussions and feedback.
Supplementary Material
Detection benchmark.
We have converted instance labels into bounding boxes for the anomaly detection benchmark. For evaluation, we considered three methods, namely UGainS [29], Mask2Anomaly [32], and VOS [11]. The COCO [26] Average Precision (AP) and Average Recall (AR) metrics serve as evaluation metrics. Unfortunately, we observed an unexpectedly poor performance of VOS. While performing well on ambiguous objects, i.e. the toy car is correctly predicted as an anomaly, vos struggles to predict for an unknown object (see Figure 4). Note that, we have not contacted the authors of VOS for help with the submission and cannot fully trust our results. We plan to open the detection benchmark for submission along with the instance benchmark, such that we can evaluate anomaly detection methods with the help of the community.
Qualitative Results.
We provide additional qualitative results in Figure 5.
Competition and Benchmark Website.
We follow a setup common [1] for hosting the benchmark. We host competition webpage (see Figure 6) on https://codalab.lisn.upsaclay.fr/ servers, and a benchmark webpage on our local server, with manually updated leaderboard for methods with at least an arXiv paper (see Figure 7).
Figure 4: VOS prediction on the Lost and Found dataset.
(a) Label
(b) U3HS
(c) Mask2Anomaly
(d) UGainS
Figure 5: Qualitative results on FS L&F, RoadAnomaly21 and RoadObstacle21 dataset.
Figure 6: Competition website overview.
Figure 7: Leaderboard on the website.

### v2

@@ IV Experiments
UGainS [26] 11.14 16.75 14.98 39.07 41.45 12.36
Mask2Anomaly [25] 01.24 02.23 00.09 18.87 19.45 09.74
U3HS [24] 00.16 00.40 01.08 01.82 01.83 03.80
LabelU3HSMask2AnomalyUGainS Fishyscapes L&F Road Anomaly21 Road Obstacle21 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Qualitative comparison of all three methods on all three datasets.
Additional Evaluation Metrics. In Tab. IV, we provide additional evaluation metrics for anomalous object detection. All performance metrics are averaged over the three datasets. The AR1, A10 and AR100 scores additionally indicate that it is particularly challenging for all the tested methods to detect the anomalous objects present in the data. The majority of objects remain overlooked, also for the most strongly performing methods, namely UGainS. Noteworthily, the small difference between AR10 and AR100 shows that all methods can hardly deal with the very crowded scenes of the dataset such as flocks of sheep on the roads.
TABLE V: Evaluation of three existing anomaly instance segmentation methods for three different buckets of object sizes in terms of AP and AP50. Higher scores indicate stronger performance.
Object Size UGainS [26] M2A [25] U3HS [24]
AP AP50 AP AP50 AP AP50
$<1,\!000$ pix 8.50 19.94 1.32 2.68 0.15 0.43
$1,\!000$-$10,\!000$ pix 34.43 55.32 30.29 53.24 0.08 0.31
$>10,\!000$ pix 17.41 31.69 32.19 48.25 0.05 0.15
Mean 25.19 42.81 13.73 24.30 0.19 0.58
Evaluation for Different Object Sizes. In Tab. V we present instance segmentation results of the three methods, averaged over all three datasets for three different size ranges. This evaluation reveals that small objects are particularly difficult to find for all three methods. From a method perspective, it is noteworthy that UGainS and Mask2Anomaly are on par in medium sized object from 1.000–10.000 pixels, and Mask2Anomaly is even slightly superior for large objects. However, it gets clearly outperformed on the small objects below 1.000 pixels, which constitute the most challenging cases of the benchmark. Note that all mean values in the benchmark are computed as weighted averages where the weighting treats all images across all datasets uniformly. Hence, the size-related results do not average out to the provided mean results.
Qualitative Results. Figure 3 provides a qualitative comparison of the three methods on all three datasets. Across all three datasets it can be observed that UH3S is able to spot the anomalous objects, however it is unable to provide accurate segmentation masks, which leads to the low numbers on the benchmarks. Mask2Anomaly already provides strong segmentation performance, however, it can be seen from the qualitative examples that some instances are overlooked or multiple instances are merged into a joint segment. UGainS is able to detect and accurately segment many of anomalous objects. However, an overproduction of false positives is observed. In summary, we conclude that there is still plenty of room do develop stronger anomaly instance segmentation and anomalous object detection methods.
V Conclusion
@@
Detecting and accurately segmenting anomaly instances on roads is a significant challenge, requiring an understanding of ’objectness’ without direct training on specific anomaly classes. In this work, we introduced new benchmarks for anomaly instance segmentation and anomalous object detection that integrate three popular anomaly datasets with new instance-wise annotations. This is complemented with an evaluation protocol that treats large and small objects as equally important. Our unified benchmark, termed OoDIS, provides a diverse set of anomalies that vary in size, rarity and the visual conditions in which they are presented. Comprising three datasets, OoDIS constitutes a challenging setting with a substantial number of images and annotation detail. We evaluate the performance of current methods for segmenting anomaly instances and provide an intuition behind the results. Our results show that current techniques struggle particularly with distant and small objects, and with precise segmentation masks. The benchmark results suggest strong opportunities for advancement in the area. As autonomous vehicle technologies continue to evolve, driven by large amounts of data, it remains a challenge to capture all possible real-world situations. Our work addresses the need to evaluate instance segmentation and object detection as a step towards reliable autonomous driving and robot navigation.
@@ Acknowledgment
We acknowledge fruitful discussions with H. Blum, R. Chan, S. Gasperini and S. Rai, as well as a contribution of annotations for Fishyscapes by H. Blum, and help with the benchmark submission from S. Gasperini and S. Rai. M.R. & M.A. acknowledge support by the German Federal Ministry of Education and Research (BMBF) within the junior research group project “UnrEAL” (grant no. 01IS22069). A.N. acknowledges funding by the BMBF project “WestAI” (grant no. 01IS22094D). We thank M. Burdorf, G. Lydakis, C. Schmidt, and others in the lab for discussions and feedback.